OpenStudy (anonymous):

A food store makes a 11–pound mixture of peanuts, almonds, and raisins. The cost of peanuts is $1.50 per pound, almonds cost $3.00 per pound, and raisins cost $1.50 per pound. The mixture calls for twice as many peanuts as almonds. The total cost of the mixture is $21.00. How much of each ingredient did the store use?

OpenStudy (anonymous):

amounts: peanuts=p, almonds=a, raisins=r,p+a+r=11,(p=2a)then 3a+r=11, 3/2p+3a+3/2r=21,(p=2a)then 4a+r=14, these two equations give us a=3, r=2 and p=6.